OpenStudy (anonymous):

The number of milligrams of a drug that remains in a patient�s system after t hours is given by the function A(t) = Iert. Juan was given 500 milligrams of medicine which leaves his bloodstream at a rate of 20%. How much of the medicine remains in his system after 6 hours?

OpenStudy (anonymous):

please define lert

OpenStudy (anonymous):

sorry it is le^rt

hero (hero):

oops

OpenStudy (anonymous):

but how do i solve that????

OpenStudy (anonymous):

total 500 ml was pumped and answer will come roughly 1600 ml remaining

hero (hero):

That's not correct. It's supposed to be less remaining, not more

OpenStudy (anonymous):

i dont think so, how can there be more than the initial amount?

OpenStudy (anonymous):

because you have not told units in le^rt , units will make the difference

OpenStudy (anonymous):

plz check units

OpenStudy (anonymous):

Its either 150.60mg 133.90mg 111.58mg 99.20mg

OpenStudy (anonymous):

i think it should be le^(-rt)

OpenStudy (anonymous):

then 150.6 will be answer

OpenStudy (anonymous):

no its le^rt

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

500/3.32=150/6

OpenStudy (anonymous):

e^1.2=3.32

OpenStudy (anonymous):

so 150.6 is the right answer

OpenStudy (anonymous):

@Hero what do you think?

OpenStudy (anonymous):

equation should be le^-rt

OpenStudy (anonymous):

150.6 is correct

OpenStudy (anonymous):

some printing error in your book

hero (hero):

Yes, @uzmashahid15 has it

hero (hero):

Since the 20% is being removed it should be -.20 not .20 so my bad

OpenStudy (anonymous):

alright thanks so much guys!!! sorry @uzmashahid15